Explain all equation of motion for constant acceleration.
1
Expert's answer
2020-10-02T07:13:37-0400

For constant acceleration, we have three equations of motion to determine variables of motion. They are-

  • Equation between initial velocity(u), final velocity(v), acceleration (a) and time(t).

v = u + at

  • Equation between displacement (s), initial velocity (u), time(t) and acceleration (a).

s = ut + "\frac{1}{2}" at2

  • Equation between final velocity (v), initial velocity (u), acceleration (a) and displacement (s).

v2-u2 = 2as